OpenStudy (anonymous):

A circle is described by the equation x2 + y2 + 14x + 2y + 14 = 0. What are the coordinates for the center of the circle and the length of the radius? (-7, -1), 36 units (7, 1), 36 units (7, 1), 6 units (-7, -1), 6 units

OpenStudy (freckles):

\[x^2+14x+(\frac{14}{2})^2=(x+\frac{14}{2})^2 \text{ or after simplifying } (x+7)^2 \\ \text{ you try this part } \\ y^2+2y+(?)^2=(x+?)^2\] what should the question mark be ?

OpenStudy (anonymous):

7

OpenStudy (freckles):

how did you come to 7?

OpenStudy (anonymous):

cut 14 in half

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

OpenStudy (freckles):

well you don't have 14 in front of y you have a 2 in front of y so why are you doing anything with 14?

OpenStudy (anonymous):

right sorry

OpenStudy (freckles):

I actually did the x part for you

OpenStudy (freckles):

I'm looking for you to try the y part using the x part I did for you as an example

OpenStudy (freckles):

but in general if you have y^2+by+c (you know if the coefficent of y^2 is 1) you just do the following: take the number in front of y....divide it by 2 and the square the result and that is what you need to add to complete the square

OpenStudy (anonymous):

so its 1

OpenStudy (freckles):

right since the number in front of our y is 2 and 2/2 is 1 and then 1^2 is 1

OpenStudy (freckles):

\[y^2+2y+1=(y+1)^2 \]

OpenStudy (freckles):

now what ever you add in to one side you need to add into the other

OpenStudy (freckles):

\[x^2+14x+y^2+2y=-14\] this is the original problem

OpenStudy (freckles):

you added in 7^2 and 1^2 on the left hand side you add need to do the same to the right hand side

OpenStudy (freckles):

\[(x+7)^2+(y+1)^2=-14+7^2+1^2\]

OpenStudy (freckles):

when you are ready you can simplify the left hand side

OpenStudy (freckles):

\[(x-h)^2+(y-k)^2=r^2 \text{ tells us } (h,k) \text{ is center and } r \text{ is radius }\]

OpenStudy (anonymous):

so it d

OpenStudy (freckles):

that is right because you get r^2=36 and so r=6 and yes (h,k)=(-7,-1)

OpenStudy (anonymous):

thank a lots

OpenStudy (anonymous):

gtg